OpenStudy (anonymous):

use the formula v1 x v2= |v1||v2| cos theta to find the cosine of the angle between the two vectors (-7, 8) and (1,0).

OpenStudy (jdoe0001):

\(\bf \textit{angle between two vectors }\\ \quad \\ cos(\theta)=\cfrac{u \cdot v}{||u||\ ||v||} \implies \cfrac{\text{dot product}}{\text{product of magnitudes}}\\ \quad \\ \theta = cos^{-1}\left(\cfrac{u \cdot v}{||u||\ ||v||}\right)\)